Calculate the new boiling point of a solution containing 5.27 g of sucrose (C12H22O11) and 45 g of water

1
Expert's answer
2022-02-19T06:55:26-0500

Solution:

A solution will boil at a higher temperature than the pure solvent.

This is the colligative property called boiling point elevation.

An equation has been developed for this behaviour. It is:

Δt = i × Kb × m

where

Δt = the change in boiling point (the boiling point of pure water is 100°C)

i = Vant Hoff's coefficient (for nonelectrolytes, like C12H22O11, i = 1)

Kb = the molal boiling-point elevation constant (Kb for water is 0.512°C kg mol¯1)

m = the molality of the solute


Molality of C12H22O11 = Moles of C12H22O11 / Kilograms solvent

Moles of C12H22O11 = Mass of C12H22O11 / Molar mass of C12H22O11

The molar mass of C12H22O11 is 342.3 g mol¯1.

Therefore,

Moles of C12H22O11 = (5.27 g) / (342.3 g mol¯1) = 0.0154 mol

Molality of C12H22O11 = (0.0154 mol) / (0.045 kg) = 0.342 mol kg¯1


Thus:

Δt = i × Kb × m = (1) × (0.512°C kg mol¯1) × (0.342 mol kg¯1) = 0.175°C = 0.18°C

 

The boiling point of pure water is 100°C.

Therefore,

t = 100°C + 0.18°C = 100.18°C

t = 100.18°C


Answer: The new boiling point of a solution is 100.18°C